Two lenses of power -15D and +15D are in contact each other.the focal lengthof the combination is.
a)+20cm
b)-10cm
c)-20cm
d)+20cm

Grade:

2 Answers

Saurabh Koranglekar
3 years ago
Dear students,
Power of combination of lens is
P1 + P2=2D + (-1.5 D) = 0.5 D

As power is positive lens is convex lens of power 0.5 D
First Lens Power = +2.D
Focal length = 1m /P = 100 cm/2 = 50 cm

focall length = 100 cm/1.5 = -67 cm
Vikas TU
3 years ago
Dear students,
Power of combination of lens is
P1 + P2=2D + (-1.5 D) = 0.5 D
As power is positive lens is convex lens of power 0.5 D
First Lens Power = +2.D
Focal length = 1m /P = 100 cm/2 = 50 cm
It is a convex lens
Ascend lens
Power = -1.5
focall length = 100 cm/1.5 = -67 cm
Itr is a concave lens with negative focal length.
Regards